# Structures of Moduli Spaces of Generalized Cantor Sets in $(0,1)^{\mathbb{N}}$

### v2

@@ 1. Introduction and main results
In a series of papers ([9], [10], [11]) , we have studied the quasiconformal equivalence of Cantor sets and their moduli spaces. For example, we show that the limit set of a Schottky group and Cantor Julia sets of hyperbolic rational functions belong to the moduli space of the standard middle one third Cantor set $\mathcal{C}_{0}$. Moreover, we show that the limit set of a generalized Schottky group with a rank one parabolic does not belong to the moduli space of $\mathcal{C}_{0}$. Recently, R. Hidalugo ([5]) shows that the number of the moduli spaces of Cantor sets generated by finitely generated Kleinian groups is at most four.
@@
We consider conditions for $\omega^{\prime}$ to belong to $\mathcal{M}(\omega)$.
@@
The result may be interesting if it is compared with the fact by R. Hidalugo mentioned above, namely, the finiteness of the number of Moduli spaces of Cantor sets generated by finitely generated Kleinian groups.
@@ 2.1. Infinite sequences in $(0,1)$ and generalized Cantor sets
For $\omega=(q_{n})_{n=1}^{\infty}\in\Omega$, we take $E_{k}(\omega)=\cup_{j=1}^{2^{k}}I_{k}^{j}$ which is the closed set at the $k$-step of the construction given in the introduction. Here, $I_{k}^{j}$ is a closed interval located at the left of $I_{k}^{j-1}$ $(j=1,2,\dots,2^{k}-1)$. The set $0,1]\setminus E_{k}(\omega)$ consists of $2^{k}-1$ open intervals $J_{k}^{1},\dots,J_{k}^{2^{k}-1}$. The interval $J_{k}^{j}$ is located between $I_{k}^{j}$ and $I_{k}^{j+1}$.
@@
For a sequence $\{\omega_{m}\}$ converging to $\omega$, we consider the complements $X(\omega_{m})$ of $E(\omega_{m})$ $(m\in\mathbb{N})$ and the Carathéodory kernel $N(\{X(\omega_{m})\})_{m=1}^{\infty}$ of the sequence of the domains. Here, $z\in\widehat{\mathbb{C}}$ belongs to the kernel $N(\{X(\omega_{m}\}_{m=1}^{\infty})$ if and only if there exist a neighborhood $U$ of $z$ and $M\in\mathbb{N}$ such that $U\subset X(\omega_{m})$ for all $m>M$.
@@
2.4. Extremal length and the modulus of a ring domain.
Let $W$ be a domain in $\widehat{\mathbb{C}}$ (or a Riemann surface) and $\Gamma$ a family of curves $\gamma$ in $W$. We consider a measurable conformal metric $\rho=\rho(z)|dz|$ on $W$ with
@@
where $\Gamma_{W}$ is the set of closed curves in $W$ separating the two connected components of $\partial W$.
A Jordan domain $Q$ together with four points located anti-clockwisely on $\partial Q$, say $x_{1},x_{2},x_{3}$ and $x_{4}$, is called a quadrilateral. We denote the quadrilateral by $(Q;x_{1},x_{2},x_{3},x_{4})$ or simply by $Q$ if it does not cause any confusion. We denote by $\overline{xy}$ for $x,y\in\partial Q$ the sub-arc of $\partial Q$ from $x$ to $y$. Then, the modulus of $(Q;x_{1},x_{2},x_{3},x_{4})$ denoted by $\textrm{Mod}(Q;x_{1},x_{2},x_{3},x_{4})$ or $\textrm{Mod}(Q)$ is defined by
$\textrm{Mod}(Q;x_{1},x_{2},x_{3},x_{4})=\textrm{Mod}(Q)=\lambda(\Gamma_{Q}),$
where $\Gamma_{Q}$ is the set of curves connecting $\overline{x_{1}x_{2}}$ and $\overline{x_{3}x_{4}}$ in $Q$.
It is known that both $\lambda(\Gamma_{W})$ and $\lambda(\Gamma_{Q})$ above have extremal metrics which attain the supremums of the definition of the extremal length (cf. [1]).
@@
Lemma 2.1.
$\lambda(\Gamma_{A(1,R)})=2\lambda(\Gamma_{A(1,R)\cap\mathbb{H}})$.
Proof.
We have
$\lambda(\Gamma_{A(1,R)})=\frac{L_{\Gamma_{A(1,r)}}(\rho_{0})^{2}}{A(\rho_{0})}$
and
[…]
Since
[…]
we obtain the desired result. ∎
@@ 3. Proof of Theorem I
Take $\omega_{0}\in\Omega$ and we show that $F_{\omega}$ is lower semi-continuous at $\omega_{0}$. Since $F_{\omega}(\omega_{0})=+\infty$ if $\omega_{0}\not\in\mathcal{M}(\omega)$, we may assume that $\omega_{0}\in\mathcal{M}(\omega)$.
For $\omega_{0}=(q_{n}^{0})_{n=1}^{\infty}\in\mathcal{M}(\omega)$, we consider $\omega_{m}=(q_{n}^{m})_{n=1}^{\infty}$ converging to $\omega_{0}$ as $m\to\infty$.
@@
$F_{\omega}(\omega_{0})\leq K(\varphi)\leq K=\liminf_{m\to\infty}F_{\omega}(\omega_{m})$
and we verify that $F_{\omega}$ is lower semi-continuous at $\omega_{0}$ as desired. ∎
@@ 4.1. The shift map and quasiconformal equivalence
The shift map $\sigma$ is a classical object in the ergodic theory. Actually, the followings are well-known.
@@
    The action of $\sigma:\Omega\to\Omega$ is ergodic. Namely, if a measurable set $A$ in $\Omega$ is invariant under the action of $\sigma$, then $m_{\infty}(A)=0$ or $1$.
@@
Here, we use the following fact on the gluing of quasiconformal mappings.
@@
Hence, $\sigma^{n}(\widehat{\omega})=\omega^{\prime}$ is quasiconformally equivalent to $\omega$ and we conclude that $\widehat{\omega}\in\mathcal{M}(\sigma^{-n}(\omega))$ from Lemma 4.1. ∎
4.2. Proof of the theorem
It is not hard to see that for a subset $S\subset\Omega$, $\sigma(\sigma^{-1}(S))=S$ and $\sigma^{-1}(\sigma(S))\supset S$, but $\sigma^{-1}(\sigma(S))\not=S$ in general. However, it is valid for moduli spaces.
Lemma 4.2.
For any $n\in\mathbb{N}$,
Proof.
It suffices to show that $\sigma^{-n}(\sigma^{n}(\mathcal{M}(\omega)))\subset\mathcal{M}(\omega)$.
@@ 4.2. Proof of the theorem
Hence, we have $\sigma^{n}(\widehat{\omega})=\sigma^{n}(\omega^{\prime})$. Thus, we conclude that $\widehat{\omega}$ is quasiconformally equivalent to $\omega^{\prime}\in\mathcal{M}(\omega)$. So, $\widehat{\omega}$ belongs to $\mathcal{M}(\omega)$ and we have $\sigma^{-n}(\sigma^{n}(\mathcal{M}(\omega)))\subset\mathcal{M}(\omega)$. ∎
@@
for any $n\in\mathbb{N}$. Also, we have from (4.3)
$m_{\infty}(\sigma^{n}(\mathcal{M}(\omega)))=m_{\infty}(\mathcal{M}(\omega)),$
@@
Hence, if $m_{\infty}(\mathcal{M}(\omega))>0$, then
@@
holds for all $n\in\mathbb{N}$ except at most finitely many $n$. Since $\mathcal{M}(\sigma^{-n}(\omega))\supset\sigma^{-n}(\mathcal{M}(\omega))$ from (4.2), we have
@@
Thus, we conclude $\mathcal{M}(\sigma^{-n}(\omega))=\mathcal{M}(\omega)$ because moduli spaces are equivalence classes. This means $\mathcal{M}(\omega)=\mathcal{M}(\sigma^{-n}(\omega))\supset\sigma^{-n}(\mathcal{M}(\omega))$. Hence, we see that $\mathcal{M}(\omega)=\sigma^{-n}(\mathcal{M}(\omega))$ as measurable sets. Therefore, $m_{\infty}(\mathcal{M}(\omega))=1$ from the ergodicity of $\sigma$ and we conclude that at most one moduli space has the positive volume, if any.
@@ 5. Proof of Theorem III
In particular, they are bounded above and converge to zero as $n\to\infty$ because $\lim_{n\to\infty}q_{n}=\lim_{n\to\infty}q^{\prime}_{n}=1$. From the hyperbolic trigonometry, we see that short geodesics do not pass through any pair of pants in the canonical pants decomposition.
@@
Proof.
It is known that a pair of pants is decomposed into congruent right-angled hexagons. Then, we see that an application of the argument of the proof of Theorem 7.19.2 in [2] §7.19 confirms the statement. ∎
@@
We consider a four-punctured sphere $X_{k,j}:=\widehat{\mathbb{C}}\setminus\cup_{i=1}^{4}x_{i}$ and we denote by $\Gamma_{0}$ the set of closed curves on $X_{k,j}$ which are homotopic to $\gamma_{k}^{j}$ in $X_{k,j}$. For each $\gamma_{0}\in\Gamma_{0}$, $\bar{\gamma}$ is the image of $\gamma$ by the reflection with respect to the real line and $\overline{\Gamma}_{0}=\{\bar{\gamma}\mid\gamma\in\Gamma_{0}\}$. We also put $\gamma^{+}=\{\gamma\cup(\overline{\gamma\cap\mathbb{L}})\}\setminus(\gamma\cap\mathbb{L})$ and $\Gamma_{0}^{+}=\{\gamma^{+}\mid\gamma\in\Gamma_{0}\}$, where $\mathbb{L}$ is the lower half plane in $\mathbb{C}$. Since $\Gamma_{0}=\overline{\Gamma_{0}}$, it follows from a symmetric principle that
$\lambda(\Gamma_{0})=\frac{1}{2}\lambda(\Gamma_{0}^{+}),$
where $\lambda(\Gamma)$ stands for the extremal length of a curve family $\Gamma$.
We denote by $\widetilde{\Gamma}_{0}$ the set of curves in the upper half plane $\mathbb{H}$ connecting two intervals $(x_{1},x_{2})$ and $(x_{3},x_{4})$. For each $\gamma\in\Gamma_{0}^{+}$, there exists $\tilde{\gamma}\in\widetilde{\Gamma}_{0}$ such that $\tilde{\gamma}\subset\gamma$. Therefore, we see
$\lambda(\widetilde{\Gamma}_{0})\leq\lambda(\Gamma_{0}^{+})$
from the definition of the extremal length.
Furthermore, from the symmetric property, we have
(5.4) $\lambda(\widehat{\Gamma}_{0})=2\lambda(\widetilde{\Gamma}_{0}),$
where $\widehat{\Gamma}_{0}$ is the set of closed curves separating $(-\infty,x_{1}]\cup[x_{4},+\infty)\cup\{\infty\}$ from $x_{2},x_{3}]$. Indeed, it is known that there exists a conformal mapping $f:\mathbb{H}\to A(1,R)\cap\mathbb{H}$ for some $R>1$ such that $f(x_{1})=-R,f(x_{2})=-1,f(x_{3})=1$ and $f(x_{4})=R$, where $A(1,R)=\{z\in\mathbb{C}\mid 1<|z|<R\}$. Since both $\mathbb{H}$ and $A(1,R)\cap\mathbb{H}$ are Jordan domains, it follows from Carathéodory’s theorem that $f$ is extended to the boundaries as a homeomorphism. Then, we define a mapping $F$ from a ring domain $A_{k,j}:=\widehat{\mathbb{C}}\setminus((-\infty,x_{1}]\cup[x_{4},+\infty)\cup\{\infty\}\cup[x_{2},x_{3}])$ onto $A(1,R)$ by
$F(z)=\begin{cases}f(z),\quad z\in\mathbb{H}\cup(-x_{1},x_{2})\cup(x_{3},x_{4})\\ \overline{f(\bar{z})},\quad z\in\mathbb{L}.\end{cases}$
From the reflection principle, the mapping $F$ is a conformal mapping from $A_{k,j}$ onto $A(1,R)$ with $F((\mathbb{H},x_{1},x_{2},x_{3},x_{4}))=(A(1,R)\cap\mathbb{H},-R,-1,1,R))$. Since the extremal length is a conformal invariant, we have
$\lambda(\widehat{\Gamma}_{0})=\lambda(\Gamma_{A(1,R)}),\quad\lambda(\widetilde{\Gamma}_{0})=\lambda(\Gamma_{A(1,R)\cap\mathbb{H}})$
and we obtain (5.4) from Lemma 2.1.
Thus, we conclude
(5.5) $\lambda(\Gamma_{0})\geq\frac{1}{4}\lambda(\widehat{\Gamma}_{0}).$
On the other hand,
(5.6) $\lambda(\widehat{\Gamma}_{0})=(\textrm{Mod}(A_{k,j}))^{-1}.$
We also see that the ring domain $A_{k,j}$ is conformally equivalent to a ring domain $B_{k,j}:=\widehat{\mathbb{C}}\setminus([-\infty,-\alpha]\cup[0,1]\cup[\beta,+\infty])$, where $\alpha=|J_{k}^{j-1}|/|I_{k}^{j}|\leq 2/(1-q_{k})$ and $\beta=1+|J_{k}^{j}|/|I_{k}.j|\leq 1+2(2/(1-q_{k}))^{k}$. Therefore,
@@
Furthermore, by taking a Möbius transformation, we see that the domain is conformally equivalent to a Teichmüller domain $A_{j,k}^{\prime\prime}:=\mathbb{C}\setminus([-1,0]\cup[\alpha^{\prime}(\beta^{\prime}-1)/(\alpha^{\prime}+\beta^{\prime}),+\infty)$ and
@@
$\lambda(\Gamma_{0})\geq\frac{1}{4}\lambda(\widehat{\Gamma}_{0})=\left(4\textrm{Mod}(A_{k,j})\right)^{-1}\geq\frac{\pi/2}{6\log 2-\log(1-q_{k})}.$
@@
On the other hand, we have seen that
@@
$\ell_{X(\omega)}(\gamma_{k}^{j})\geq 2e^{-C(\delta)}\lambda(\Gamma_{0})\geq\frac{e^{-C(\delta)}\pi}{6\log 2-\log(1-q_{k})}.$
@@
$\ell_{X(\omega^{\prime})}({\gamma^{\prime}}_{k}^{j})\geq\frac{e^{-C(\delta)}\pi}{6\log 2-\log(1-q^{\prime}_{k})}..$
@@
(5.8) $\frac{e^{-C(\delta)}\pi}{6\log 2-\log(1-q_{k})}\leq\ell_{X(\omega)}(\gamma_{k}^{j})\leq\frac{2\pi^{2}}{\log\left(1+\frac{2\delta}{1-q_{k}}\right)},$
@@
(5.9) $\frac{e^{-C(\delta)}\pi}{6\log 2-\log(1-q^{\prime}_{k})}\leq\ell_{X(\omega)}({\gamma^{\prime}}_{k}^{j})\leq\frac{2\pi^{2}}{\log\left(1+\frac{2\delta}{1-q_{k}^{\prime}}\right)}$
@@
$R(q_{k},q_{k}^{\prime})$ $:=$ $\frac{2\pi e^{C(\delta)}(6\log 2-\log(1-q^{\prime}_{k}))}{\log\left(1+\frac{2\delta}{1-q_{k}}\right)}$
$=$ $2\pi e^{C(\delta)}\frac{\log(1-q_{k}^{\prime})}{\log(1-q_{k})}\cdot\frac{\frac{6\log 2}{\log(1-q_{k}^{\prime})}-1}{\frac{\log(2\delta+1-q_{k})}{\log(1-q_{k})}-1}.$
@@
$R(q_{k}^{\prime},q_{k})$ $:=$ $\frac{2\pi e^{C(\delta)}(6\log 2-\log(1-q_{k}))}{\log\left(1+\frac{2\delta}{1-q_{k}^{\prime}}\right)}$
$=$ $2\pi e^{C(\delta)}\frac{\log(1-q_{k})}{\log(1-q_{k}^{\prime})}\cdot\frac{\frac{6\log 2}{\log(1-q_{k}^{\prime})}-1}{\frac{\log(2\delta+1-q_{k}^{\prime})}{\log(1-q_{k}^{\prime})}-1}.$
@@
Since we are supposing (5.1), there exists a subsequence $\{n_{k}\}_{k=1}^{\infty}$ such that $\lim_{k\to\infty}q_{n_{k}}=\lim_{k\to\infty}q^{\prime}_{n_{k}}=1$. In particular, $\lim_{k\to\infty}\ell({\gamma^{\prime}}_{n_{k}}^{j})=0$ by (5.9).
@@
Step 3 : Comparing the lengths of geodesics.
@@
$=$ $KR(q^{\prime}_{n_{k}},q_{n_{k}})\cdot\frac{e^{C(\delta)}\pi}{6\log 2-\log(1-q_{k})}$
@@
$\ell_{X(\omega)}([\varphi({\gamma^{\prime}}_{n_{k}}^{j})])$ $\leq$ $KR(q^{\prime}_{n_{k}},q_{n_{k}})\cdot\frac{e^{C(\delta)}\pi}{6\log 2-\log\left(1-q_{n_{k}}\right)}$
$\leq$ $KR(q^{\prime}_{n_{k}},q_{n_{k}})\cdot\frac{e^{C(\delta)}\pi}{6\log 2+\log\left(1-q_{n}\right)}<\ell_{X(\omega)}(\gamma_{n}^{m})$

### v3

@@ 1. Introduction and main results
In a series of papers ([11], [12], [13]) , we have studied the quasiconformal equivalence of Cantor sets and their moduli spaces. For example, we show that the limit set of a Schottky group and Cantor Julia sets of hyperbolic rational functions belong to the moduli space of the standard middle one third Cantor set $\mathcal{C}_{0}$. Moreover, we show that the limit set of a generalized Schottky group with a rank one parabolic does not belong to the moduli space of $\mathcal{C}_{0}$. Recently, R. Hidalgo ([7]) shows that the number of the moduli spaces of Cantor sets generated by finitely generated Kleinian groups is at most four.
@@
In [13], we have conjectured that $m_{\infty}(\mathcal{M}(\omega))=0$ for any $\omega\in\Omega$. While we have given a partial affirmative answer in Theorem II, the conjecture is still open.
We consider conditions for $\omega^{\prime}\in\Omega$ to belong to $\mathcal{M}(\omega)$.
@@
The result may be interesting if it is compared with the fact by R. Hidalgo mentioned above, namely, the finiteness of the number of Moduli spaces of Cantor sets generated by finitely generated Kleinian groups.
Acknowledgment. The author thanks the anonymous referee for helpful comments and suggestions which improve this paper.
@@ 2.1. Infinite sequences in $(0,1)$ and generalized Cantor sets
For $\omega=(q_{n})_{n=1}^{\infty}\in\Omega$, we take $E_{k}(\omega)=\cup_{j=1}^{2^{k}}I_{k}^{j}$ which is the closed set at the $k$-step of the construction given in the introduction. Here, $I_{k}^{j}$ is a closed interval located at the right of $I_{k}^{j-1}$ $(j=1,2,\dots,2^{k}-1)$. The set $0,1]\setminus E_{k}(\omega)$ consists of $2^{k}-1$ open intervals $J_{k}^{1},\dots,J_{k}^{2^{k}-1}$. The interval $J_{k}^{j}$ is located between $I_{k}^{j}$ and $I_{k}^{j+1}$.
@@
For a sequence $\{\omega_{m}\}$ converging to $\omega$, we consider the complements $X(\omega_{m})$ of $E(\omega_{m})$ $(m\in\mathbb{N})$ and the Carathéodory kernel $N(\{X(\omega_{m})\}_{m=1}^{\infty})$ of the sequence of the domains. Here, $z\in\widehat{\mathbb{C}}$ belongs to the kernel $N(\{X(\omega_{m}\}_{m=1}^{\infty})$ if and only if there exist a neighborhood $U$ of $z$ and $M\in\mathbb{N}$ such that $U\subset X(\omega_{m})$ for all $m>M$ (cf. [9]).
@@
2.4. Extremal length, the modulus of a ring domain and holomorphic quadratic differentials.
Let $W$ be a domain in $\widehat{\mathbb{C}}$ (or a Riemann surface) and $\Gamma$ a family of curves in $W$. We consider a measurable conformal metric $\rho=\rho(z)|dz|$ on $W$ with
@@
where $\Gamma_{W}$ is the set of closed curves in $W$ separating the two connected components of $\partial W$. It is known that both $\lambda(\Gamma_{W})$ and $\lambda(\Gamma_{Q})$ above have extremal metrics which attain the supremums of the definition of the extremal length (cf. [1]).
@@
Let $R$ be a hyperbolic Riemann surface of finite type and $\varphi=\varphi(z)dz^{2}$ be a holomorphic quadratic differential on $R$ with $\|\varphi\|:=\iint|\varphi|<\infty$. Then, a branch of $w=\int_{p_{0}}^{p}\sqrt{\varphi}$ gives a local coordinate around $p_{0}\in R$ unless $p_{0}$ is a zero of $\varphi$. A horizontal trajectory of $\varphi$ is the locus of $\{\textrm{Re }w=c\}$ for some $c\in\mathbb{R}$. We say that a horizontal trajectory of $\varphi$ is closed if it is a simple closed curve on $R$. Then, the following is known:
Proposition 2.2.
Let $\gamma$ be a simple closed hyperbolic geodesic on $R$. Then, there exists a holomorphic quadratic differential $\varphi_{\gamma}$ on $R$ with $\|\varphi_{\gamma}\|<\infty$ such that
  1. (1)
    any closed horizontal trajectory of $\varphi_{\gamma}$ is homotopic to $\gamma$,
  2. (2)
    the set of closed horizontal trajectories of $\varphi_{\gamma}$ forms an annulus $A_{\gamma}$ and
    $\textrm{Mod}(A_{\gamma})=\lambda(\Gamma_{\gamma})^{-1},$
    where $\Gamma_{\gamma}$ is the set of simple closed curves homotopic to $\gamma$ on $R$.
@@ 3. Proof of Theorem I
Take $\omega_{0}\in\Omega$ and we show that $F_{\omega}$ is lower semi-continuous at $\omega_{0}$.
At first, we assume that $\omega_{0}\in\mathcal{M}(\omega)$. For $\omega_{0}=(q_{n}^{0})_{n=1}^{\infty}\in\mathcal{M}(\omega)$, we consider $\omega_{m}=(q_{n}^{m})_{n=1}^{\infty}$ converging to $\omega_{0}$ as $m\to\infty$.
@@
$F_{\omega}(\omega_{0})\leq K(\varphi)\leq K=\liminf_{m\to\infty}F_{\omega}(\omega_{m}).$
If $\omega_{0}\not\in\mathcal{M}(\omega)$, then $F_{\omega}(\omega_{0})=+\infty$. Take a sequence $\{\omega_{m}\}_{m=1}^{\infty}$ in $(0,1)^{\mathbb{N}}$ which converges to $\omega_{0}$.
If $\omega_{m}\not\in\mathcal{M}(\omega)$ for any $m\in\mathbb{N}$, then $F_{\omega}(\omega_{m})=+\infty$ and we have
$F_{\omega}(\omega_{0})=+\infty=\lim_{m\to\infty}F_{\omega}(\omega_{m}).$
If the sequence contains infinitely many elements in $\mathcal{M}(\omega)$, say $\omega_{m_{k}}$ $(k=1,2,\dots)$, then we obtain
$\lim_{k\to\infty}F_{\omega}(\omega_{m_{k}})=+\infty=F_{\omega}(\omega_{0}).$
Indeed, if $\liminf_{k\to\infty}F_{\omega}(\omega_{m_{k}})<+\infty$, then we conclude that $F_{\omega}(\omega_{0})<+\infty$ by using the normal family argument as in the first part of the proof above and it is a contradiction.
Thus, we verify that $F_{\omega}$ is lower semi-continuous at $\omega_{0}$ as desired. ∎
@@ 4.1. The shift map and quasiconformal equivalence
The shift map $\sigma$ is a classical object in the ergodic theory. Actually, the followings are well-known (cf. [5]).
@@
    The action of $\sigma:\Omega\to\Omega$ is mixing with respect to $m_{\infty}$. Namely, for any measurable sets $A,B$ in $\Omega$,
    (4.1) $\lim_{n\to\infty}m_{\infty}(A\cap\sigma^{-n}(B))=m_{\infty}(A)m_{\infty}(B)$
    holds. As mixing is a stronger property than ergodicity, the action of $\sigma$ on $\Omega$ is ergodic.
@@
Here, we use the following fact on the gluing of quasiconformal mappings ( [11] Lemma 4.1) .
@@
Hence, $\sigma^{n}(\widehat{\omega})=\omega^{\prime}$ is quasiconformally equivalent to $\omega$ and we conclude that $\widehat{\omega}\in\mathcal{M}(\sigma^{-n}(\omega))$ from Lemma 4.1.
As for (4.4), it suffices to show that $\sigma^{-n}(\sigma^{n}(\mathcal{M}(\omega)))\subset\mathcal{M}(\omega)$ because $\sigma^{-n}(\sigma^{n}(\mathcal{M}(\omega)))\supset\mathcal{M}(\omega)$ is obvious.
@@
Hence, we have $\sigma^{n}(\widehat{\omega})=\sigma^{n}(\omega^{\prime})$. Thus, we conclude that $\widehat{\omega}$ is quasiconformally equivalent to $\omega^{\prime}\in\mathcal{M}(\omega)$. So, $\widehat{\omega}$ belongs to $\mathcal{M}(\omega)$ and we have $\sigma^{-n}(\sigma^{n}(\mathcal{M}(\omega)))\subset\mathcal{M}(\omega)$.
4.2. Proof of the theorem
@@
Suppose that $m_{\infty}(\mathcal{M}(\omega))>0$. Then, from the mixing property (4.1) we see that there exists $N\in\mathbb{N}$ such that
@@
Since $\mathcal{M}(\sigma^{-N}(\omega))\supset\sigma^{-N}(\mathcal{M}(\omega))$ from (4.3), we have
@@
Thus, we conclude $\mathcal{M}(\sigma^{-N}(\omega))=\mathcal{M}(\omega)$ because moduli spaces are equivalence classes. This means $\mathcal{M}(\omega)=\mathcal{M}(\sigma^{-N}(\omega))\supset\sigma^{-N}(\mathcal{M}(\omega))$. Hence, we see that $\mathcal{M}(\omega)=\sigma^{-N}(\mathcal{M}(\omega))$ as measurable sets. Since $\sigma^{N}$ is also mixing, it follows from (4.1) that $m_{\infty}(\mathcal{M}(\omega))=1$ if $m_{\infty}(\mathcal{M}(\omega))>0$. Thus, we conclude that at most one moduli space has the positive volume, if any.
@@ 5. Proof of Theorem III
In particular, they are bounded above and converge to zero as $n\to\infty$ because $\lim_{n\to\infty}q_{n}=\lim_{n\to\infty}q^{\prime}_{n}=1$. From the hyperbolic trigonometry, we see that short geodesics do not pass through any pair of pants in the canonical pants decomposition. Actually, from the collar lemma (cf. [4] Chapter 4), we may obtain the following lemma immediately.
@@
We consider a four-punctured sphere $X_{k,j}:=\widehat{\mathbb{C}}\setminus\cup_{i=1}^{4}x_{i}$ and we denote by $\Gamma_{0}$ the set of closed curves on $X_{k,j}$ which are homotopic to $\gamma_{k}^{j}$ in $X_{k,j}$. Now, we consider a holomorphic quadratic differential $\varphi_{0}$ on $X_{k,j}$ defined by
$\varphi_{0}:=\frac{dz^{2}}{(z-x_{1})(z-x_{2})(z-x_{3})(z-x_{4})}.$
We see that any closed horizontal trajectory of $\varphi_{0}$ belongs to $\Gamma_{0}$. We also see that the set of closed horizontal trajectories of $\varphi_{0}$ forms a ring domain $A_{k,j}:=\widehat{\mathbb{C}}\setminus((-\infty,x_{1}]\cup[x_{4},+\infty)\cup\{\infty\}\cup[x_{2},x_{3}])$. Hence, from Proposition 2.2, we have
(5.4) $\lambda(\Gamma_{0})=(\textrm{Mod}(A_{k,j}))^{-1}.$
It is seen that the ring domain $A_{k,j}$ is conformally equivalent to a ring domain $B_{k,j}:=\widehat{\mathbb{C}}\setminus([-\infty,-\alpha]\cup[0,1]\cup[\beta,+\infty])$, where $\alpha=|J_{k}^{j-1}|/|I_{k}^{j}|\leq 2/(1-q_{k})$ and $\beta=1+|J_{k}^{j}|/|I_{k}.j|\leq 1+2(2/(1-q_{k}))^{k}$. Therefore,
@@
Furthermore, by taking a Möbius transformation, we see that the domain $A_{k,j}^{\prime}$ is conformally equivalent to a Teichmüller domain $A_{j,k}^{\prime\prime}:=\mathbb{C}\setminus([-1,0]\cup[\alpha^{\prime}(\beta^{\prime}-1)/(\alpha^{\prime}+\beta^{\prime}),+\infty)$ and
@@
$\lambda(\Gamma_{0})=\left(\textrm{Mod}(A_{k,j})\right)^{-1}\geq\frac{1}{6\log 2-\log(1-q_{k})}.$
@@
On the other hand, we have seen in (5.2) that
@@
$\ell_{X(\omega)}(\gamma_{k}^{j})\geq 2e^{-C(\delta)}\lambda(\Gamma_{0})\geq\frac{2e^{-C(\delta)}}{6\log 2-\log(1-q_{k})}.$
@@
$\ell_{X(\omega^{\prime})}({\gamma^{\prime}}_{k}^{j})\geq\frac{2e^{-C(\delta)}}{6\log 2-\log(1-q^{\prime}_{k})}..$
@@
(5.6) $\frac{2e^{-C(\delta)}}{6\log 2-\log(1-q_{k})}\leq\ell_{X(\omega)}(\gamma_{k}^{j})\leq\frac{2\pi^{2}}{\log\left(1+\frac{2\delta}{1-q_{k}}\right)},$
@@
(5.7) $\frac{2e^{-C(\delta)}}{6\log 2-\log(1-q^{\prime}_{k})}\leq\ell_{X(\omega)}({\gamma^{\prime}}_{k}^{j})\leq\frac{2\pi^{2}}{\log\left(1+\frac{2\delta}{1-q_{k}^{\prime}}\right)}$
@@
$R(q_{k},q_{k}^{\prime})$ $:=$ $\frac{\pi^{2}e^{C(\delta)}(6\log 2-\log(1-q^{\prime}_{k}))}{\log\left(1+\frac{2\delta}{1-q_{k}}\right)}$
$=$ $\pi^{2}e^{C(\delta)}\frac{\log(1-q_{k}^{\prime})}{\log(1-q_{k})}\cdot\frac{\frac{6\log 2}{\log(1-q_{k}^{\prime})}-1}{\frac{\log(2\delta+1-q_{k})}{\log(1-q_{k})}-1}.$
@@
$R(q_{k}^{\prime},q_{k})$ $:=$ $\frac{\pi^{2}e^{C(\delta)}(6\log 2-\log(1-q_{k}))}{\log\left(1+\frac{2\delta}{1-q_{k}^{\prime}}\right)}$
$=$ $\pi^{2}e^{C(\delta)}\frac{\log(1-q_{k})}{\log(1-q_{k}^{\prime})}\cdot\frac{\frac{6\log 2}{\log(1-q_{k}^{\prime})}-1}{\frac{\log(2\delta+1-q_{k}^{\prime})}{\log(1-q_{k}^{\prime})}-1}.$
@@
Since we are supposing (5.1), there exists a subsequence $\{n_{k}\}_{k=1}^{\infty}$ such that $\lim_{k\to\infty}q_{n_{k}}=\lim_{k\to\infty}q^{\prime}_{n_{k}}=1$ and
(5.8) $\lim_{k\to\infty}\left|\log\frac{\log(1-q_{n_{k}})}{\log(1-q_{n_{k}^{\prime}})}\right|=\infty.$
In particular, $\lim_{k\to\infty}\ell({\gamma^{\prime}}_{n_{k}}^{j})=0$ by (5.7).
@@
Step 4 : Comparing the lengths of geodesics.
@@
$=$ $KR(q^{\prime}_{n_{k}},q_{n_{k}})\cdot\frac{2e^{C(\delta)}}{6\log 2-\log(1-q_{k})}$
@@
$\ell_{X(\omega)}([\varphi({\gamma^{\prime}}_{n_{k}}^{j})])$ $\leq$ $KR(q^{\prime}_{n_{k}},q_{n_{k}})\cdot\frac{2e^{C(\delta)}}{6\log 2-\log\left(1-q_{n_{k}}\right)}$
$\leq$ $KR(q^{\prime}_{n_{k}},q_{n_{k}})\cdot\frac{2e^{C(\delta)}}{6\log 2-\log\left(1-q_{n}\right)}<\ell_{X(\omega)}(\gamma_{n}^{m})$